OpenStudy (jekdidkdjrjrjjr):

FAN+MEDAL SUPER EASY 2x + 3y = -17 is one equation in a coincidental system of two linear equations. The other equation is_______________ y =_____________ x – 102.

OpenStudy (jekdidkdjrjrjjr):

still need help

OpenStudy (mathmale):

First of all, what does "coincidental" mean in this context? Have you looked up this word?

OpenStudy (jekdidkdjrjrjjr):

@mathmale yeah but i still don’t get it

OpenStudy (mathmale):

First of all, what does "coincidental" mean in this context?

OpenStudy (jekdidkdjrjrjjr):

@calculusxy

OpenStudy (jekdidkdjrjrjjr):

@mathmale happening by or resulting from coincidence? lmao

OpenStudy (mathmale):

What does this word mean in the context of algebra / geometry?

OpenStudy (jekdidkdjrjrjjr):

@TheSmartOne

OpenStudy (calculusxy):

Coincidental system of two linear equations, when graphed, will be placed on the same line--in other words, they have infinite common points.

OpenStudy (calculusxy):

You need another equation that will be equivalent to 2x + 3y = -17

OpenStudy (calculusxy):

One strategy that I usually utilize during these situations to get an equivalent equation is that I multiply all of the terms by a certain number. In this case, we can multiply it by 2. \(2(2x + 3y = -17)\) which would equal to \(4x + 6y = - 34\) You would have two linear equations that are coincidental: \(2x + 3y = -17\) \(4x + 6y = -34\) To check, you can graph them.